OpenStudy (anonymous):

how would you simplify: (n+1)! (n^n)/((n+1)^(n+1))(n!)

OpenStudy (anonymous):

\[(n+1)!n ^{n}/(n+1)^{n+1}n!\]

OpenStudy (zarkon):

\[\frac{(n+1)!}{n!}=n+1\] \[\frac{n^n}{(n+1)^{n+1}}=\frac{n^n}{(n+1)^n(n+1)}=\left(\frac{n}{n+1}\right)^n\frac{1}{n+1}\]

OpenStudy (anonymous):

so what is the final answer?